OpenStudy (calculusxy):

Rodney has fifty coins, including at least one quarter, that are worth $1.00. If he loses one coin, what is the probability that it was a dime?

OpenStudy (calculusxy):

@phi

OpenStudy (phi):

let me know when you figure it out.

OpenStudy (calculusxy):

I have no idea on how to figure this out.

OpenStudy (phi):

you have to examine the possibilities. they say you have at least one quarter. say you had 2 quarters and 48 other coins (assumed to be pennies, nickels or dimes) and they add up to 100 cents if we subtract the 2 quarters off the 100 cents, the 48 coins have to add up to 50 cents Let p, n and d be the number of pennies nickels and dimes for this case, you have p+n+d = 48 (48 coins) from which we get p= 48-n-d also, the value of the coins is p +5n +10d = 50 cents replace p with 48-n-d and simplify 48 +4n + 9d = 50 4n + 9n = 2 if we have even 1 nickel or 1 dime, we will get a number bigger than 2 all of this means we cannot have 2 or more quarters.

OpenStudy (anonymous):

the only way to make $1 with 50 coins is, 1Q, 2D, 2N, 45P. So 2/50

OpenStudy (phi):

because there is at least 1 quarter, we have exactly 1 quarter, and 49 other coins we have p+n+d= 49 which means p= 49 - n - d p+ 5n+10d= 75 cents (the quarter is worth the remainder 25 cents) replace p with 49-n - d and simplify 4n + 9 d = 26

OpenStudy (phi):

now you have 4n + 9 d = 26 if you have 3 dimes, you get 27... which is too big. so you know you have 0, 1 or 2 dimes. test the 3 choices and see if you can get an *integer* number of nickels